OpenStudy (anonymous):

Tickets for a school play cost $5 if you buy them early, and $7 if you buy them the day of the show. If 200 tickets are sold, and the total amount of money from ticket sales was $1340, how many tickets were purchased early? A. 170 tickets B. 30 tickets C. 100 tickets D. 124 tickets

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@dietrich_harmon how? ;o

OpenStudy (anonymous):

No kidding.

OpenStudy (anonymous):

Let early tickets be \(x\). Since there are 200 total, late tickets must be \(200 - x\). You also know that the sum of all tickets is $1340. So: \[5x + 7(200-x) = 1340\] Does that make sense to you?

OpenStudy (anonymous):

@geoffb So it really is C?

OpenStudy (anonymous):

Definitely not.

OpenStudy (anonymous):

I thought so @geoffb. I thought it was A.

OpenStudy (anonymous):

Careful...

OpenStudy (anonymous):

Did you work out the formula above? Does it make sense how I got it and what it means?

OpenStudy (anonymous):

So Id end up with 4x+1400=1340?

OpenStudy (anonymous):

No, not 4x.

OpenStudy (anonymous):

\[5x + 1400 - 7x = 1340\]

OpenStudy (anonymous):

-2x+1400=1340

OpenStudy (anonymous):

Yes, good.

OpenStudy (anonymous):

x=30

OpenStudy (anonymous):

Very good! So once you solve for x, a very important step is to go back and look at what you chose x to represent. In this case, we chose for it to represent tickets purchased early. If we had let it represent tickets purchased late, you would end up with x = 170, but the answer *wouldn't* be A, because it is looking for tickets purchased early, not late.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

OH so its B

OpenStudy (anonymous):

Yes, and you can test it if you want (I would if you were doing a test. 30 early tickets x $5 = $150 170 late tickets x $7 = $1190 Total ticket revenue = $1340

OpenStudy (anonymous):

thanks @geoffb

OpenStudy (anonymous):

You're welcome! :)